In a class of 8 boys and 15 girls four children were absent. What fraction of the children were a) absent. b)present.
Answer by checkley77(12844)

8+15=23 TOTAL CHILDREN
23-4=19 CHILDREN PRESENT.
19/23=.826=82.6 PERCENT PRESENT.
4/23=.174=17.4 PERCENT ABSENT.